SHOW ALL CALCULATIONS

You have recently accepted a position with Vitex, Inc., the manufacturer of a popular consumer product. During your first week on the job, the vice president has been favorably impressed with your work. She has been so impressed, in fact, that yesterday she called you into her office and asked you to attend the executive committee meeting this morning for the purpose of leading a discussion on the variances reported for last period. Anxious to favorably impress the executive committee, you took the variances and supporting data home last night to study On your way to work this morning, the papers were laying on the seat of your new, red convertible. As you were crossing a bridge on the highway, a sudden gust of wind caught the papers and blew them over the edge of the bridge and into the stream below. You managed to retrieve only one page, which contains the following information Standard Cost Card Direct materials, 2.40 pounds at $16.40 per pound Direct labor, 1.00 direct labor-hours at $15.20 per direct labor-hour Variable manufacturing overhead, 1.00 direct labor-hours at S9.30 per direct labor-hour $ 39.36 15.20 $ 9.30 Variances Reported Direct materials Direct labor Variable manufacturing overhead Total Standard Cost* $629,760 $ 243,200 $148,800 Price or Rate $11,716 F $ 3,400 U $4,300 F Quantity or Efficiency $32,800 U $15,200 U You recall that manufacturing overhead cost is applied to production on the basis of direct labor-hours and that all of the materials purchased during the period were used in production. Work in process inventories are insignificant and can be ignored It is now 8:30 a.m. The executive committee meeting starts in just one hour, you realize that to avoid looking like a bungling fool you must somehow generate the necessary backup data for the variances before the meeting begins. Without backup data it will be impossible to lead the discussion or answer any questions

How many units were produced last period?

How many pounds of direct material were purchased and used in production?

What was the actual cost per pound of material?

How many actual direct labor-hours were worked during the period?

What was the actual rate paid per direct labor-hour?

How much actual variable manufacturing overhead cost was incurred during the period?
Answer #1

Particular

Rate/Unit

Cost

Direct Material

39.36

626,760

Direct Labour

15.20

243,200

Variable Manufacturing overhead

9.30

148,800

Standard Quantity = Total Cost/Rate per unit

                                     =   626,760/39.36

                                    =   15,923.78 units

Standard Hours = Total Rate/ Rate per hour

                                = 243,200/15.20

                                = 16,000 hours

Material Quantity Variance = (Standard Quantity – Actual Quantity) * Standard Price per unit

-32,800 =    (15,923.78- AQ)*39.36

Actual Quantity= 32800/39.36+15,923.78

     = 16757.11 Units

Material Price Variance = (Standard Price per unit – Actual Price per unit) * Actual Quantity

11,716 = (39.36-AP)*16,757.11

Actual Price = -11,716/16,757.11 +    39.36

                       = $38.66

Direct Labour Efficiency Variance = (Standard hours – Actual hours) * standard rate per hour

-15,200 =   (16,000- Actual hours) * 15.2

Actual Hours = 15,200/15.2 +   16,000

                           = 17,000 hours

Direct Labour Price Variance= (Standard rate per hour – actual rate per hour) * Actual Hours

-3,400    =             (15.2- Actual rate per hour)*17,000

Actual rate per hour = 3400/17000 + 15.2

                                         = $15.4

Variable manufacturing overhead price variance= (Standard rate per hour – actual rate per hour) *                  Actual Hours

4,300 = (9.20- Actual rate per hour)*17,000         

Variable manufacturing actual rate per hour = -4,300/17,000   + 9.20

                                                                                        = $ 8.94

Variable Manufacturing overhead Efficiency Variance = (Standard hours – Actual hours) * standard rate per hour

                                                                                     = (16,000-17,000) *9.3

                                                                                             = $9,300 U

Answers:

How Many units were produced last periods = 16,757.11 units

How many pounds of direct material were purchased and used in production:

Pounds required for 1 quantity = 2.40 pounds

So for producing 16,757.11 total pounds require = 16,757.11*2.4

                                                                                                = 40,217.06 pounds

What was the actual cost per pound of material?   : $38.66

How many actual direct labour-hours were worked during the period? = 17,000 hours

What was the actual rate paid per direct labour-hour? = $15.40

How much actual variable manufacturing overhead cost was incurred during the period?

Actual hours * Actual rate per hour

17,000 * $8.94

$151,980